### v1

# QUIC Delay Control: an implementation of congestion and delay control

Abstract.
A new congestion and delay control algorithm named QUIC Delay Control (QUIC-DC) is proposed for controlling not only congestion but also the queuing delay encountered along the forward communication path. The core idea is to estimate the one-way queuing delay of a connection to trigger an early reaction to congestion. This idea, along with the TCP Westwood+ congestion control algorithm, has been implemented in QUIC-DC and compared with QUIC Cubic, BBRv2, NewReno, Westwood+. The results obtained in both emulated and real network connections show that QUIC-DC can significantly reduce packet losses along with end-to-end communication delays, while preserving network utilization, features that are both very useful for real-time applications.
Congestion Control, Delay Control, QUIC
ccs: Network protocols  low-delay congestion control
1. Introduction
Classic congestion control aims at maximizing throughput while minimizing congestion, i.e. packet losses and retransmissions. This implies that communication delays are not controlled and oscillate between a minimum, equal to the propagation time, up to a maximum equal to the propagation time plus the maximum queuing delay encountered along the communication path.
The goal of this paper is to propose an algorithm that not only avoids congestion while still providing high network utilization but also controls the network queuing delay.
Controlling delays is nowadays of the utmost importance since the Internet is not only aimed at delay insensitive data traffic, but is increasingly used for real-time applications such as live streaming, video-conferencing, virtual reality/augmented reality, $360^{\circ}$ video, autonomous driving, tele-robotics and tele-surgery.
An Internet connection is established using an end-to-end transport protocol, the TCP or the UDP, over the connectionless Internet Protocol (IP) to allow a bidirectional communication between a pair of sender/receiver nodes identified by two IP addresses.
Figure 1 shows a connection between two TCP/UDP peers over an IP network. The essential features that play a role and need to be considered are the communication links and the buffers. Indeed, TCP/IP or UDP/IP packets go through a set of communication links and buffers connecting the routers encountered as they travel along the communication path from the sender to the destination. Figure 1 shows the sequence of communication links and buffers, each link $i$ with a propagation delay $T_{p_{i}}$ and each buffer $i$ with a queuing delay $T_{q_{i}}$, encountered by a packet/datagram traveling from the sender to the destination.
The round trip time ($RTT$) is the sum of the propagation and queuing delays encountered going from the sender to the destination and then back to the sender: $RTT=\sum_{i\in P}(T_{p_{i}}+T_{q_{i}})$, where $P$ is the set of links and buffers belonging to the bidirectional communication path.
The propagation delays $T_{p_{i}}$ are fixed and depend on the physical distance (the propagation speed is roughly 1/3 of the speed of light). In contrast, the queuing delays $T_{q_{i}}$ depend on queuing occupancy and can be reduced by controlling the queue length. The overall queuing delay is $T_{q}=\sum_{i\in P}T_{q_{i}}$, whereas the overall propagation delay is $T_{p}=\sum_{i\in P}T_{p_{i}}$.
In other terms, the RTT is made of a constant propagation time plus a time-varying stochastic component due to the queuing delay as follows: $RTT=T_{p}+T_{q}$
QUIC-DC can bound the component of delay due to the queuing. Conceptually, it can be viewed as a form of Explicit Congestion Notification (ECN) (Ramakrishnan et al., 2001) implemented end-to-end.
Figure 1. Packets going from the sender to the receiver through a series of links and buffers
2. Related Work
Traditional loss-based TCP (Jacobson, 1988; Peterson and Davie, 2021) is not suitable for delay-sensitive traffic, such as in the case of video conferencing traffic, because its congestion control mechanism continuously probes for available network bandwidth, creating a periodic pattern in which network queues are cyclically filled up to packet loss and then drained. These queue oscillations introduce a time-varying stochastic queuing delay component, which adds to the fixed network propagation delay, thereby impairing delay-sensitive communications. But more than anything, it is the retransmission of lost packets that harms real-time communication. Therefore reducing the packet loss rate is fundamental in real-time traffic. For this reason, the idea of using delay measurements to proactively react to the onset of network congestion has been proposed (Jain, 1989). Issues related to delay measurements and throughput degradation when loss-based flows share the bottleneck with delay-based flows have been addressed in (Grieco and Mascolo, 2004; Budzisz et al., 2011; Prasad et al., 2004). The first efforts focusing on the reduction of the queuing delay were set in the TCP congestion control research domain. In particular, the first congestion control algorithm specifically designed to contain the end-to-end latency by employing delay measurements is described in the seminal work by Jain (Jain, 1989). Since then, many delay-based TCP congestion control variants have been designed, such as TCP Vegas (Brakmo and Peterson, 1995) or  (Budzisz et al., 2011). TCP Vegas employs end-to-end RTT measurements to establish delay thresholds used to infer incipient congestion by comparison with a base RTT (Brakmo and Peterson, 1995). It has been shown that, when the RTT is used as a congestion metric, a low channel utilization may be obtained in the presence of reverse traffic, which inflates queues in the backward path, or when competing with loss-based flows (Grieco and Mascolo, 2004). To tackle such fairness issues, the algorithms proposed in (Budzisz et al., 2011) infer the presence of concurrent loss-based flows and switch to loss-based mode, returning back to the delay-based mode when loss-based flows are detected to be no longer present. Another class of algorithms employs one-way-delay measurements to rule out the sensitivity to reverse-path congestion. Examples are LEDBAT (over UDP) (Shalunov et al., 2012) and TCP Santa Cruz (Parsa and Garcia-Luna-Aceves, 1999). In particular, LEDBAT employs an increasing phase in which the congestion window is increased at a rate that is proportional to the distance between the measured one-way delay and a fixed delay target (Shalunov et al., 2012). It has been shown that LEDBAT is affected by the so-called latecomer effect: when two flows share the same bottleneck the second flow typically starves the first one (Carofiglio et al., 2010). Recently the idea of employing RTT gradient has been proposed to overcome the aforementioned latecomer effect. Some examples are CDG (Hayes and Armitage, 2011) and Verus (Zaki et al., 2015). CDG has been designed with the aim of coexisting with loss-based flows while keeping end-to-end delay low (Hayes and Armitage, 2011). Verus has been designed for the case of cellular networks where sudden link capacity variations make the congestion control design challenging (Zaki et al., 2015). To the best of our knowledge the only congestion control algorithm for real-time flows that has been implemented and massively deployed is the Google Congestion Control (GCC) (Carlucci et al., 2017). The algorithm has been implemented in the Google Chrome web browser. GCC employs the one way delay variation as a congestion signal and an adaptive threshold mechanism to effectively compete with loss-based congestion control algorithms (Carlucci et al., 2017).
3. QUIC Delay Control
In this section, we introduce the main idea to enforce queuing delay control. The algorithm will be implemented and evaluated in QUIC (Iyengar and Thomson, 2021).
To focus on the delay problem, let us consider again Figure 1, which shows a connection between two TCP/UDP peers over an IP network. As we have seen, the RTT is the sum of the encountered propagation delays and queuing delays.
Therefore, by keeping the queue as low as possible, we can target a minimum RTT equal to the propagation delay only:
(1) $\ {RTT}_{\text{min}}\ =T_{p}$
Similarly, $\text{RTT}_{\text{max}}$ can be defined as:
(2) ${RTT}_{\text{max}}={RTT}_{\text{min}}+T_{q_{\max}}$
where $T_{q_{\max}}$ is the maximum queuing time encountered by a packet going through all the buffers belonging to the communication path.
Figure 2. Measurement of one-way delay variation
At this point, it is important to distinguish between queuing on the forward path and queuing on the backward path. A connection is primarily constrained by congestion on the forward path; therefore, if RTT is used as a congestion signal, queuing in the backward path caused by reverse traffic may lead to a false indication of congestion. This, in turn, would unnecessarily slow down the forward traffic — a behavior that is indeed observed in TCP Vegas (Grieco and Mascolo, 2004).
To this end, it is necessary to consider the variation of the One-Way Delay (OWD), where the OWD is the time a packet takes to travel from the sender to the receiver, in order to estimate the forward queuing delay only. Precisely, the OWD is the difference between the packet arrival time $t^{r}$ at the receiver and the packet departure time $t^{s}$ at the sender as shown in Figure 2:
(3) $OWD_{i}=t^{r}_{i}-t^{s}_{i}$
The OWD measurement contains a systematic error due to the clock offset between the sender and the receiver. However, what is useful to know is not the absolute value but just the OWD variation $OWDV$, which measures an increase or decrease in the forward queuing delay. Thus, by computing the one-way delay variation:
(4) $\begin{split}&OWDV_{i}={OWD_{i}-OWD_{i-1}}\\ &=(t^{r}_{i}-t^{s}_{i})-(t^{r}_{i-1}-t^{s}_{i-1})\\ &=(t^{r}_{i}-t^{r}_{i-1})-(t^{s}_{i}-t^{s}_{i-1})\end{split}$
Equation 4 shows that the one-way delay variation can be easily obtained by computing the difference between the inter arrival times of two consecutive packets $i$ and $i$$1$, and their corresponding inter departure times.
By adding the one way delay variations, we can compute the one-way queuing delay as:
(5) $OWQD_{i}=OWQD_{i-1}+OWDV_{i}$
It should be noted that this measure is both simple and robust and can be easily implemented.
Once we have the measurement of the queuing delay along the forward path OWQD, we can use it to define a new congestion event which occurs when $OWQD$ reaches a threshold $OWQD_{th}$:
(6) $OWQD>OWQD_{th}$
Using Eq. (6), we can control queue build-up and reduce the risk of packet loss, which is particularly important for latency-sensitive applications.
In QUIC Delay Control (QUIC-DC), we use the TCP Westwood+ congestion control (Grieco and Mascolo, 2004; Mascolo et al., 2001) i.e., after a congestion episode, QUIC-DC sets the congestion window equal to the product of the available bandwidth times the minimum $RTT$, which keeps full the propagation pipe and empty the queuing pipe.
In summary, a congestion episode is defined by a timeout, a 3DUPACK, or $\text{OWD}>\text{OWD}_{th}$. The reaction to a congestion episode is the same as that of TCP Westwood after a timeout or after 3DUPACK, and the reaction to $\text{OWD}>\text{OWD}_{th}$ is the same of TCP Westwood after 3DUPACK.
The implementation of TCP Westwood+ congestion control (Grieco and Mascolo, 2004) in QUIC-DC closely adheres to its counterpart implementation in the Linux TCP kernel with a small change in the Westwood low-pass filter to make the bandwidth estimation BWE faster. Indeed, the original low pass filter used to estimate the available bandwidth BWE adversely impacts the algorithm’s responsiveness during transient network conditions, leading to a lower goodput during the start up phase. The low pass filter employed in QUIC-DC is:
(7) $BWE_{i}=0.2\cdot BWE_{i-1}+0.8\cdot bandwidthsample_{i}$
which significantly improves the goodput.
4. QUIC-DC implementation and the testbed
We have implemented and integrated the Westwood+ congestion control algorithm and the new QUIC Delay Control into Meta’s mvfst QUIC transport implementation, which serves as the underlying transport layer for Proxygen, Meta’s HTTP libraries collection [^1] https://engineering.fb.com/2014/11/05/production-engineering/introducing-proxygen-facebook-s-c-http-framework. Both mvfst and Proxygen are implemented in C++ to ensure high performance and scalability. To simplify deployment and testing, all components were containerized using Docker, providing a consistent and reproducible environment. In our testbed, we have set the maximum possible value for the advertised window so that the congestion window size is constrained only by the network capacity.
The testbed for the evaluation of QUIC-DC comprises two main architectures. The first is a controlled laboratory testbed made of two machines connected with an Ethernet cable. The first machine acts as the client, sending one or multiple HTTP3 / QUIC requests to the second one operates as the server. Both machines are equipped with Gigabit NICs. All the Docker containers are based on Ubuntu 22.04 images. In this case, the bottleneck is located physically on the interface of the server where we applied a 25 ms delay both incoming and outgoing, thus summing up to a total propagation time of 50 ms. We use the Linux traffic control (tc) tool from the iproute2 suite to perform all the network emulation tasks [^2] https://man7.org/linux/man-pages/man8/tc.8.html. The propagation delay $T_{p}$ is set using netem queuing disciplines on the server Ethernet interface, while bandwidth capacity $C$ in Mbps and buffer size in bytes are set using the token bucket filter (TBF) mechanism. The results are averaged over 3 runs and are consistent between each run.
The first uncontrolled real Internet setup consists of a residential client in Bari that accesses the network using an ADSL modem. The server is located at the University of Rome, La Sapienza.
Figure 3. Satellite testbed scenario.
The second real network testbed is made of users accessing Starlink and connected to other Internet users as in Figure 3. The connection was established between nodes at Politecnico di Bari, Italy and Karlstad University, Sweden.
Since the experimental results are affected by the uncontrolled coexisting network conditions, considered QUIC congestion control algorithms are executed in sequence and the sequence is repeated 40 times. Pacing is enabled at the application level at the server for all the experiments to ensure that packets are transmitted at evenly spaced intervals, minimizing burstiness.  [^3] The pacing interval has been set to 200 $\mu s$.
5. Experimental results
This section presents the experimental results that have been run to validate QUIC-DC. To the purpose, QUIC-DC has been compared with other state-of-the-art congestion control algorithms, namely Westwood+ (Grieco and Mascolo, 2004), Cubic (Ha et al., 2008), New Reno (Henderson et al., 2012), and BBR version 2  (Cardwell et al., 2019). Notice that, even though BBRv3 is available, in this work we consider BBRv2 due to its superior fairness compared to BBRv3 (Zeynali et al., 2024). The bottleneck link capacity has been set equal to $C$ Mbps and the bottleneck buffer size has been set equal 0.5, 1, 2, 4 times the bandwidth delay product (BDP) where the bandwidth is the bottleneck capacity $C$ and the delay is the round trip propagation delay $T_{p}=RTT_{min}$.
In the following, we show results in the cases of single-flow and multiple-flow connections.
5.1. Emulated network results
5.1.1. Single Flow
To understand the dynamics, we begin by examining a single-flow connection transferring a 100 MB file. QUIC-DC has been tested using different values for the one-way queuing delay threshold, denoted as $OWQD_{th}$ set to 10%, 20%, 50%, and 80% of the bottleneck buffer size expressed in time.
Figure 4. Round‑trip‑time (RTT) distribution for a single flow – emulated link, buffer=4×BDP
Table 1 shows the performance metrics of a single flow connections for the case of a buffer size set equal to 4 BDP, where $C$=10 Mbps and $RTT_{min}=T_{p}=$50 ms. In this case, QUIC-DC reaches maximum goodput (Gput in the table) with minimum loss percentage compared to the other congestion control algorithms independently of the considered queuing delay threshold $OWQD_{th}$ and lower RTT. In comparison to QUIC Cubic, the loss percentage is reduced by 2 orders of magnitude! Figure 4 shows the corresponding CDF of the measured RTT for all the algorithms under analysis. The curves show that QUIC-DC provides lower RTT values compared to the other algorithms. Moreover, it can be noticed that by decreasing the $OWQD_{th}$ the RTT curves shift to the left towards lower RTT values. This is expected since a higher $OWQD_{th}$ implies larger queues along the path. It is worth noting that in the case of algorithms with a larger loss rate, the RTT measurements underestimate the effective delays because retransmitted packets experience very large delays, which do not contribute to the RTT measurements (Karn and Partridge, 1987). Similar considerations can be drawn from Table 2, which corresponds to the case of bottleneck buffer size set to 2 BDP.
A nice understanding of the proposed approach is provided by Figure 5, which compares the $RTT$ and $cwnd$ dynamics of QUIC Westwood+ (blue line) and QUIC-DC (red line) in the case of a bottleneck buffer size set to 2 BDP. It can be observed that, while QUIC Westwood+ resets the congestion window when the maximum RTT is reached (dashed green line), QUIC-DC resets the congestion window when the $OWQD$ reaches the delay control threshold set at 80% of the bottleneck buffer size in time. Under this setting, buffer saturation at the bottleneck is prevented, effectively avoiding packet losses. Consequently, the observed packet loss ratio was reduced by one order of magnitude (see Table 1).
When the buffer size is set to 1 BDP or 0.5 BDP, results reported in Table 3 and 4 show that QUIC-DC tends to be less efficient, in terms of goodput. It is worth noting that, in the case of congestion control algorithms that experience higher packet loss—such as BBR2—the measurement of packet RTT does not account for retransmitted packets (Karn and Partridge, 1987). As a result, the RTT values tend to underestimate the actual delays.
Figure 5. Time evolution of RTT and congestion window – buffer=2×BDP.
Table 1. Single-flow emulation; buffer = $4\times\text{BDP}$ ( $C=10\,\text{Mbps}$, $RTT_{\min}=50\,\text{ms}$ ).
CCA Gput Tput Loss RTTavg RTTstd
(Mbps) (Mbps) (%) (ms) (ms)
QUIC-DC (10%) 9.16 9.16 0.02 55.75 3.00
QUIC-DC (20%) 9.17 9.17 0.01 63.83 7.64
QUIC-DC (50%) 9.16 9.17 0.03 88.60 19.14
QUIC-DC (80%) 9.16 9.16 0.02 123.82 26.48
Westwood+ 9.18 9.19 0.07 170.23 43.10
BBRv2 9.15 9.21 0.65 102.33 27.44
Cubic 9.10 9.20 1.03 190.32 31.37
New Reno 9.20 9.20 0.07 159.26 45.14
Table 2. Single-flow emulation; buffer = $2\times\text{BDP}$ ( $C=10\,\text{Mbps}$, $RTT_{\min}=50\,\text{ms}$ ).
CCA Gput Tput Loss RTTavg RTTstd
(Mbps) (Mbps) (%) (ms) (ms)
QUIC-DC (10%) 9.06 9.06 0.03 52.74 1.23
QUIC-DC (20%) 9.16 9.16 0.03 55.98 3.65
QUIC-DC (50%) 9.14 9.15 0.14 68.33 10.25
QUIC-DC (80%) 9.16 9.16 0.04 84.86 16.47
Westwood+ 9.15 9.18 0.26 90.53 20.04
BBRv2 9.12 9.18 0.66 87.94 26.94
Cubic 9.16 9.20 0.39 104.82 13.57
New Reno 9.17 9.18 0.14 93.78 18.79
Table 3. Single-flow emulation; buffer = $1\times\text{BDP}$ ( $C=10\,\text{Mbps}$, $RTT_{\min}=50\,\text{ms}$ ).
CCA Gput Tput Loss RTTavg RTTstd
(Mbps) (Mbps) (%) (ms) (ms)
QUIC-DC (10%) 7.08 7.13 0.63 52.51 1.06
QUIC-DC (20%) 7.75 7.79 0.53 52.63 1.65
QUIC-DC (50%) 8.51 8.58 0.83 53.64 3.96
QUIC-DC (80%) 8.39 8.46 0.83 52.44 1.15
Westwood+ 8.22 8.27 0.64 53.11 3.03
BBRv2 9.12 9.21 0.87 57.21 7.72
Cubic 9.19 9.20 0.15 63.06 7.15
New Reno 7.39 7.42 0.37 52.39 1.10
Table 4. Single-flow emulation; buffer = $0.5\times\text{BDP}$ ( $C=10\,\text{Mbps}$, $RTT_{\min}=50\,\text{ms}$ ).
CCA Gput Tput Loss RTTavg RTTstd
(Mbps) (Mbps) (%) (ms) (ms)
QUIC-DC (10%) 3.30 3.32 0.44 52.45 1.49
QUIC-DC (20%) 3.58 3.62 1.24 52.23 1.06
QUIC-DC (50%) 3.44 3.49 1.42 52.11 1.02
QUIC-DC (80%) 3.75 3.82 1.77 52.26 1.04
Westwood+ 3.68 3.73 1.33 52.29 1.02
BBRv2 6.27 6.35 1.20 51.41 0.77
Cubic 5.53 5.54 0.24 51.32 0.95
New Reno 3.34 3.37 0.81 52.48 0.96
5.1.2. Multiple Flows
Let us now consider the multi-flow scenario shown in Figure 3, where four connections of the same type share the same channel and are initiated at different start times, i.e. they enter the channel every 10 seconds, and they leave it in a staggered way as well. In this case, the control delay threshold $OWQD_{th}$ is set to 80% for QUIC-DC, and the bottleneck buffer size is set to 2 BDP.
Table 5. Four concurrent flows – emulated link, buffer=2×BDP. (Average $\pm$ standard deviation).
CCA RTTavg Goodput Loss
(ms) (Mbps, JFI) (%)
QUIC-DC (80%) $\boldsymbol{95.07}\pm\boldsymbol{0.16}$ $2.29\pm 0.07$ (0.999) $\boldsymbol{0.12}\pm\boldsymbol{0.02}$
Westwood+ $101.19\pm 0.13$ $2.28\pm 0.12$ (0.998) $0.43\pm 0.06$
BBRv2 $111.55\pm 0.29$ $2.28\pm 0.34$ (0.984) $1.05\pm 0.11$
Cubic $114.26\pm 0.09$ $2.30\pm 0.03$ (0.999) $0.34\pm 0.03$
New Reno $113.87\pm 0.44$ $1.95\pm 0.12$ (0.998) $0.58\pm 0.07$
Table 5 presents the average values along with the corresponding standard deviations for the RTT, goodput, Jain’s fairness index, and loss percentage of each control algorithm. As shown, QUIC-DC exhibits the lowest RTT and loss percentage, while maintaining a goodput level comparable to the other considered algorithms. It is worth noting that, since the link capacity is set to 10 Mbps, a goodput around 2.5 Mbps indicates that the link is fairly shared among the four connections.
Figure 6. RTT distribution for four concurrent flows – emulated link, buffer=2×BDP.
Figure 6 shows the CDF of the RTT in the case of multiple connections over a shared bottleneck. The figure shows that QUIC-DC provides lower RTT values compared to the other considered congestion control algorithms. Due to space constraints, we are unable to extend our investigation to more complex multiflows scenarios, nor can we explore the friendliness between different types of control algorithms, which are left for further investigation.
5.2. Real network results
In this section we report results obtained in the real network scenario described in Section 4.
5.2.1. Terrestrial network
During the experiments, we have measured $RTT{min}=30\ ms$, $RTT{max}=60\ ms$, which means that there is a bottleneck buffer size equal to $1\ BDP$. Table 6 shows that QUIC-DC reduces the loss rate by one order of magnitude whereas the goodput is only slightly lower than the other approaches (note that QUIC Westwood+ corresponds to QUIC-DC $(100\%)$). Moreover, Figure 7 shows that QUIC-DC also provides shorter RTTs. It should be noted that algorithms experiencing higher loss rates provide RTT delays that are underestimated because retransmitted packets are not considered in RTT measurements (Karn and Partridge, 1987).
Figure 7. RTT distribution on real terrestrial path.
Table 6. Terrestrial path, single flow.
CCA Gput Tput Loss RTTavg RTTstd
(Mbps) (Mbps) (%) (ms) (ms)
QUIC-DC (20%) 26.744 26.746 0.0096 30.92 9.30
QUIC-DC (50%) 30.954 30.956 0.0103 32.09 11.18
Westwood+ 33.115 33.132 0.0470 35.08 11.50
BBRv2 33.560 33.586 0.0730 32.00 11.28
Cubic 33.529 33.555 0.0733 36.17 12.05
New Reno 33.935 32.959 0.0765 35.92 12.18
5.2.2. Starlink LEO network
We consider the satellite scenario depicted in Figure 3. During the experiments, we measured $RTT_{min}=45\ \mathrm{ms}$ and $RTT_{max}\simeq 100\ \mathrm{ms}$, implying a queuing delay of circa 55 ms. This corresponds to a buffer size of approximately $1\times\mathrm{BDP}$. We’ve conducted experiments with a single flow and with multiple concurrent flows with the same methodology shown in Section 5.1.Table 7 shows that BBRv2 obtains a significant larger goodput, a result that is not kept in the case of multi-flows. Table 8 shows that BBRv2 is broken in the multi-flow case, where it gets 1/4 of the QUIC Westwood+ goodput with a loss rate 2.4 times higher.
Figure8 shows the CDF of the RTTs for all the considered algorithms in the case of single flow. QUIC CUBIC obtains a slightly better distribution but with a smaller goodput as shown in Table 7. Figure 9 shows the CDF of the RTTs for all the considered algorithms in the case of multiple flows. In this case QUIC BBRv2 provides the worst distribution even though it provides the worst goodput (only 5 Mbps whereas QUIC Westwood+ provides 19 Mbps). QUIC-DC using different thresholds on the one-way delay can slightly improve the RTT distributions. However, it should also be noted that algorithms experiencing higher loss rates provide RTT delays that are underestimated because retransmitted packets are not considered in RTT measurements (Karn and Partridge, 1987). In other terms, the algorithms with larger loss-rates provide a distribution of RTT that underestimates the effective delay.
Figure 8. RTT distribution on LEO satellite path – single flow
Table 7. Starlink LEO satellite path, single flow.
CCA Gput Tput Loss RTTavg RTTstd
(Mbps) (Mbps) (%) (ms) (ms)
QUIC-DC (20%) 4.81 4.82 0.13 60.58 7.51
QUIC-DC (50%) 18.06 18.08 0.09 62.30 9.41
Westwood+ 20.12 20.29 0.80 63.21 11.82
BBRv2 29.85 30.17 1.04 67.70 27.02
Cubic 15.07 15.19 0.82 59.29 18.31
New Reno 9.45 9.64 1.96 66.05 20.83
Figure 9. RTT distribution on Starlink LEO satellite path – four concurrent flows
Table 8. Starlink LEO satellite path, four concurrent flows.
CCA RTTavg Goodput Loss
(ms) (Mbps, JFI) (%)
QUIC-DC (20%) $66.36\pm 0.35$ $4.05\pm 0.07$ (1.000) $0.10\pm 0.01$
QUIC-DC (50%) $72.29\pm 0.98$ $8.70\pm 0.61$ (0.996) $0.09\pm 0.02$
Westwood+ $81.72\pm 1.27$ $19.17\pm 0.95$ (0.998) $1.02\pm 0.45$
BBRv2 $87.04\pm 10.50$ $5.09\pm 1.60$ (0.931) $2.44\pm 0.63$
Cubic $74.76\pm 2.78$ $15.28\pm 1.63$ (0.992) $0.81\pm 0.09$
New Reno $78.70\pm 2.06$ $14.81\pm 1.49$ (0.992) $0.93\pm 0.41$
6. Conclusions
In this paper, we presented QUIC-DC, an algorithm that explicitly controls the queuing delay of a communication connection. The core idea is to estimate the connection one-way queuing delay in order to trigger an early reaction to congestion. The idea has been implemented in QUIC using the TCP Westwood+ congestion control. QUIC-DC has been evaluated against QUIC NewReno, QUIC Westwood+, QUIC Cubic, and QUIC BBR2. Experimental results show that the cumulative distribution function of the round-trip time can be shifted to lower values, while the packet loss rate can be reduced by an order of magnitude.

### v3

@@
 Abstract.
-A new congestion and delay control algorithm named QUIC Delay Control (QUIC-DC) is proposed for controlling not only congestion but also the queuing delay encountered along the forward communication path.
-The core idea is to estimate the one-way queuing delay of a connection to trigger an early reaction to congestion.
-This idea, along with the TCP Westwood+ congestion control algorithm, has been implemented in QUIC-DC and compared with QUIC Cubic, BBRv2, NewReno, Westwood+.
-The results obtained in both emulated and real network connections show that QUIC-DC can significantly reduce packet losses along with end-to-end communication delays, while preserving network utilization, features that are both very useful for real-time applications.
+A new congestion and delay control algorithm named QUIC Delay Control (QUIC-DC) is proposed for controlling not only congestion but also the queueing delay encountered along the forward communication path.
+The core idea is to estimate the one-way queueing delay of a connection to trigger an early reaction to congestion.
+This idea, along with a the TCP Westwood+ congestion control algorithm, has been implemented in QUIC-DC and compared with QUIC Cubic, BBRv2, NewReno, Westwood+.
+The results obtained in the emulated and real network environments show that QUIC-DC can significantly reduce packet losses along with end-to-end communication delays, while preserving network utilization, features that are both very useful for real-time applications.
 Congestion Control, Delay Control, QUIC
-ccs: Network protocols  low-delay congestion control
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM/IRTF Applied Networking Research Workshop 2025; July 19-25, 2025; Madrid, Spainisbn: 978-1-4503-XXXX-X/2018/06ccs: Network protocols  low-delay congestion control
 1. Introduction
@@
 This implies that communication delays are not controlled and oscillate between a minimum, equal to the propagation time, up to a maximum equal to the propagation time plus the maximum queuing delay encountered along the communication path.
-The goal of this paper is to propose an algorithm that not only avoids congestion while still providing high network utilization but also controls the network queuing delay.
+The goal of this paper is to propose an algorithm that not only avoids congestion while still providing high network utilization but also controls the network queueing delay.
 Controlling delays is nowadays of the utmost importance since the Internet is not only aimed at delay insensitive data traffic, but is increasingly used for real-time applications such as live streaming, video-conferencing, virtual reality/augmented reality, $360^{\circ}$ video, autonomous driving, tele-robotics and tele-surgery.
-An Internet connection is established using an end-to-end transport protocol, the TCP or the UDP, over the connectionless Internet Protocol (IP) to allow a bidirectional communication between a pair of sender/receiver nodes identified by two IP addresses.
-Figure 1 shows a connection between two TCP/UDP peers over an IP network.
-The essential features that play a role and need to be considered are the communication links and the buffers.
-Indeed, TCP/IP or UDP/IP packets go through a set of communication links and buffers connecting the routers encountered as they travel along the communication path from the sender to the destination.
-Figure 1 shows the sequence of communication links and buffers, each link $i$ with a propagation delay $T_{p_{i}}$ and each buffer $i$ with a queuing delay $T_{q_{i}}$, encountered by a packet/datagram traveling from the sender to the destination.
-The round trip time ($RTT$) is the sum of the propagation and queuing delays encountered going from the sender to the destination and then back to the sender: $RTT=\sum_{i\in P}(T_{p_{i}}+T_{q_{i}})$, where $P$ is the set of links and buffers belonging to the bidirectional communication path.
+In this paper we propose a new congestion and delay control and we experiment it in controlled and uncontrolled network environments.
+Figure 1 shows an Internet connection between two TCP/UDP peers over an IP network.
+The essential elements that play a key role are the communication links and the buffers.
+Indeed, TCP/IP or UDP/IP packets go through a set of communication links and buffers connecting the routers that are encountered as they travel from the sender to the destination.
+Figure 1 shows the sequence of communication links and buffers, each link $i$ with a propagation delay $T_{p_{i}}$ and each buffer $i$ with a queueing delay $T_{q_{i}}$, which are encountered by a packet travelling from the sender to the destination.
+The round trip time ($RTT$) is the sum of the propagation and queueing delays encountered going from the sender to the destination and then back to the sender: $RTT=\sum_{i\in P}(T_{p_{i}}+T_{q_{i}})$, where $P$ is the set of links and buffers belonging to the bidirectional communication path.
 The propagation delays $T_{p_{i}}$ are fixed and depend on the physical distance (the propagation speed is roughly 1/3 of the speed of light).
-In contrast, the queuing delays $T_{q_{i}}$ depend on queuing occupancy and can be reduced by controlling the queue length.
-The overall queuing delay is $T_{q}=\sum_{i\in P}T_{q_{i}}$, whereas the overall propagation delay is $T_{p}=\sum_{i\in P}T_{p_{i}}$.
-In other terms, the RTT is made of a constant propagation time plus a time-varying stochastic component due to the queuing delay as follows: $RTT=T_{p}+T_{q}$
-QUIC-DC can bound the component of delay due to the queuing. Conceptually, it can be viewed as a form of Explicit Congestion Notification (ECN) (Ramakrishnan et al., 2001) implemented end-to-end.
+In contrast, the queueing delays $T_{q_{i}}$ depend on queueing occupancy and can be reduced by controlling the queue length.
+The overall queueing delay is $T_{q}=\sum_{i\in P}T_{q_{i}}$, whereas the overall propagation delay is $T_{p}=\sum_{i\in P}T_{p_{i}}$.
+In other terms, the RTT is made of a constant propagation time plus a time-varying stochastic component due to the queueing delay as follows: $RTT=T_{p}+T_{q}$
+QUIC-DC can bound the component of delay due to the queueing. Conceptually, it can be viewed as a form of Explicit Congestion Notification (ECN) (Ramakrishnan et al., 2001) implemented end-to-end.
 Figure 1. Packets going from the sender to the receiver through a series of links and buffers
@@
 Traditional loss-based TCP (Jacobson, 1988; Peterson and Davie, 2021) is not suitable for delay-sensitive traffic, such as in the case of video conferencing traffic, because its congestion control mechanism continuously probes for available network bandwidth, creating a periodic pattern in which network queues are cyclically filled up to packet loss and then drained.
-These queue oscillations introduce a time-varying stochastic queuing delay component, which adds to the fixed network propagation delay, thereby impairing delay-sensitive communications.
+These queue oscillations introduce a time-varying stochastic queueing delay component, which adds to the fixed network propagation delay, thereby impairing delay-sensitive communications.
 But more than anything, it is the retransmission of lost packets that harms real-time communication.
-Therefore reducing the packet loss rate is fundamental in real-time traffic.
+Therefore, reducing the packet loss rate is fundamental in real-time traffic.
 For this reason, the idea of using delay measurements to proactively react to the onset of network congestion has been proposed (Jain, 1989).
 Issues related to delay measurements and throughput degradation when loss-based flows share the bottleneck with delay-based flows have been addressed in (Grieco and Mascolo, 2004; Budzisz et al., 2011; Prasad et al., 2004).
-The first efforts focusing on the reduction of the queuing delay were set in the TCP congestion control research domain.
+The first efforts focusing on the reduction of the queueing delay were set in the TCP congestion control research domain.
 In particular, the first congestion control algorithm specifically designed to contain the end-to-end latency by employing delay measurements is described in the seminal work by Jain (Jain, 1989).
@@
 To tackle such fairness issues, the algorithms proposed in (Budzisz et al., 2011) infer the presence of concurrent loss-based flows and switch to loss-based mode, returning back to the delay-based mode when loss-based flows are detected to be no longer present.
-Another class of algorithms employs one-way-delay measurements to rule out the sensitivity to reverse-path congestion.
+Other algorithms employ one-way-delay measurements to rule out the sensitivity to reverse-path congestion.
 Examples are LEDBAT (over UDP) (Shalunov et al., 2012) and TCP Santa Cruz (Parsa and Garcia-Luna-Aceves, 1999).
@@
 Verus has been designed for the case of cellular networks where sudden link capacity variations make the congestion control design challenging (Zaki et al., 2015).
-To the best of our knowledge the only congestion control algorithm for real-time flows that has been implemented and massively deployed is the Google Congestion Control (GCC) (Carlucci et al., 2017).
-The algorithm has been implemented in the Google Chrome web browser.
-GCC employs the one way delay variation as a congestion signal and an adaptive threshold mechanism to effectively compete with loss-based congestion control algorithms (Carlucci et al., 2017).
+The Google Congestion Control (GCC) (Carlucci et al., 2017), implemented in the Google Chrome web browser, employs the one way delay variation as a congestion signal and an adaptive threshold mechanism to effectively compete with loss-based congestion control algorithms (Carlucci et al., 2017).
+GCC is the only widely deployed non-proprietary congestion control algorithm used in a video conferencing application such as Google Meet.
 3. QUIC Delay Control
-In this section, we introduce the main idea to enforce queuing delay control. The algorithm will be implemented and evaluated in QUIC (Iyengar and Thomson, 2021).
+In this section, we introduce the main idea to obtain queueing delay control. The algorithm will be implemented and evaluated in QUIC (Iyengar and Thomson, 2021).
 To focus on the delay problem, let us consider again Figure 1, which shows a connection between two TCP/UDP peers over an IP network.
-As we have seen, the RTT is the sum of the encountered propagation delays and queuing delays.
+As we have seen, the RTT is the sum of the encountered propagation delays and queueing delays.
 Therefore, by keeping the queue as low as possible, we can target a minimum RTT equal to the propagation delay only:
@@
 (2) ${RTT}_{\text{max}}={RTT}_{\text{min}}+T_{q_{\max}}$
-where $T_{q_{\max}}$ is the maximum queuing time encountered by a packet going through all the buffers belonging to the communication path.
+where $T_{q_{\max}}$ is the maximum queueing time encountered by a packet going through all the buffers belonging to the communication path.
 Figure 2. Measurement of one-way delay variation
-At this point, it is important to distinguish between queuing on the forward path and queuing on the backward path.
-A connection is primarily constrained by congestion on the forward path; therefore, if RTT is used as a congestion signal, queuing in the backward path caused by reverse traffic may lead to a false indication of congestion.
+At this point, it is important to distinguish between queueing on the forward path and queueing on the backward path.
+A connection is primarily constrained by congestion on the forward path; therefore, if RTT is used as a congestion signal, queueing in the backward path caused by reverse traffic may lead to a false indication of congestion.
 This, in turn, would unnecessarily slow down the forward traffic — a behavior that is indeed observed in TCP Vegas (Grieco and Mascolo, 2004).
-To this end, it is necessary to consider the variation of the One-Way Delay (OWD), where the OWD is the time a packet takes to travel from the sender to the receiver, in order to estimate the forward queuing delay only.
+To this end, it is necessary to consider the variation of the One-Way Delay (OWD), where the OWD is the time a packet takes to travel from the sender to the receiver, in order to estimate the forward queueing delay only.
 Precisely, the OWD is the difference between the packet arrival time $t^{r}$ at the receiver and the packet departure time $t^{s}$ at the sender as shown in Figure 2:
@@
 The OWD measurement contains a systematic error due to the clock offset between the sender and the receiver.
-However, what is useful to know is not the absolute value but just the OWD variation $OWDV$, which measures an increase or decrease in the forward queuing delay.
+However, what is useful to know is not the absolute value but just the OWD variation $OWDV$, which measures an increase or decrease in the forward queueing delay.
 Thus, by computing the one-way delay variation:
@@
 Equation 4 shows that the one-way delay variation can be easily obtained by computing the difference between the inter arrival times of two consecutive packets $i$ and $i$$1$, and their corresponding inter departure times.
-By adding the one way delay variations, we can compute the one-way queuing delay as:
+By summing the one way delay variations, we can compute the one-way queueing delay as:
 (5) $OWQD_{i}=OWQD_{i-1}+OWDV_{i}$
 It should be noted that this measure is both simple and robust and can be easily implemented.
-Once we have the measurement of the queuing delay along the forward path OWQD, we can use it to define a new congestion event which occurs when $OWQD$ reaches a threshold $OWQD_{th}$:
+Once we have the measurement of the queueing delay along the forward path, we can use it to define a new congestion event which occurs when $OWQD$ reaches a threshold $OWQD_{th}$:
 (6) $OWQD>OWQD_{th}$
 Using Eq. (6), we can control queue build-up and reduce the risk of packet loss, which is particularly important for latency-sensitive applications.
-In QUIC Delay Control (QUIC-DC), we use the TCP Westwood+ congestion control (Grieco and Mascolo, 2004; Mascolo et al., 2001) i.e., after a congestion episode, QUIC-DC sets the congestion window equal to the product of the available bandwidth times the minimum $RTT$, which keeps full the propagation pipe and empty the queuing pipe.
+In QUIC Delay Control (QUIC-DC), we use the TCP Westwood+ congestion control (Grieco and Mascolo, 2004; Mascolo et al., 2001) i.e., after a congestion episode, QUIC-DC sets the congestion window equal to the product of the available bandwidth times the minimum $RTT$, which keeps full the propagation pipe and empty the queueing pipe.
 In summary, a congestion episode is defined by a timeout, a 3DUPACK, or $\text{OWD}>\text{OWD}_{th}$.
-The reaction to a congestion episode is the same as that of TCP Westwood after a timeout or after 3DUPACK, and the reaction to $\text{OWD}>\text{OWD}_{th}$ is the same of TCP Westwood after 3DUPACK.
-The implementation of TCP Westwood+ congestion control (Grieco and Mascolo, 2004) in QUIC-DC closely adheres to its counterpart implementation in the Linux TCP kernel with a small change in the Westwood low-pass filter to make the bandwidth estimation BWE faster.
+The reaction to a congestion episode is the same of that of TCP Westwood+ after a timeout or after 3DUPACK, and the reaction to $\text{OWD}>\text{OWD}_{th}$ is the same of TCP Westwood+ after 3DUPACK.
+The implementation of TCP Westwood+ congestion control (Grieco and Mascolo, 2004) in QUIC-DC closely adheres to its counterpart implementation in the Linux TCP kernel with a small change in the Westwood+ low-pass filter to make the bandwidth estimation BWE faster.
 Indeed, the original low pass filter used to estimate the available bandwidth BWE adversely impacts the algorithm’s responsiveness during transient network conditions, leading to a lower goodput during the start up phase.
@@
 4. QUIC-DC implementation and the testbed
-We have implemented and integrated the Westwood+ congestion control algorithm and the new QUIC Delay Control into Meta’s mvfst QUIC transport implementation, which serves as the underlying transport layer for Proxygen, Meta’s HTTP libraries collection [^1] https://engineering.fb.com/2014/11/05/production-engineering/introducing-proxygen-facebook-s-c-http-framework.
+QUIC-DC has been implemented into Meta’s mvfst QUIC transport implementation, which serves as the underlying transport layer for Proxygen, Meta’s HTTP libraries collection [^1] https://engineering.fb.com/2014/11/05/production-engineering/introducing-proxygen-facebook-s-c-http-framework.
 Both mvfst and Proxygen are implemented in C++ to ensure high performance and scalability.
-To simplify deployment and testing, all components were containerized using Docker, providing a consistent and reproducible environment.
-In our testbed, we have set the maximum possible value for the advertised window so that the congestion window size is constrained only by the network capacity.
-The testbed for the evaluation of QUIC-DC comprises two main architectures.
+To simplify deployment and testing, all components were containerized using Docker.
+In our laboratory testbed, we have set the maximum possible value for the advertised window so that the congestion window size is constrained only by the congestion window.
+QUIC-DC has been evaluated using two testbeds.
 The first is a controlled laboratory testbed made of two machines connected with an Ethernet cable.
-The first machine acts as the client, sending one or multiple HTTP3 / QUIC requests to the second one operates as the server.
-Both machines are equipped with Gigabit NICs.
+The first machine acts as the client, sending one or multiple HTTP3/QUIC requests to the second machine, which acts as the server.
+Both machines are equipped with Gigabit Ethernet Network Interface Controllers (NICs).
 All the Docker containers are based on Ubuntu 22.04 images.
-In this case, the bottleneck is located physically on the interface of the server where we applied a 25 ms delay both incoming and outgoing, thus summing up to a total propagation time of 50 ms.
+The bottleneck is located physically on the interface of the server where we applied a 25 ms delay both incoming and outgoing, thus summing up to a total propagation time of 50 ms.
 We use the Linux traffic control (tc) tool from the iproute2 suite to perform all the network emulation tasks [^2] https://man7.org/linux/man-pages/man8/tc.8.html.
-The propagation delay $T_{p}$ is set using netem queuing disciplines on the server Ethernet interface, while bandwidth capacity $C$ in Mbps and buffer size in bytes are set using the token bucket filter (TBF) mechanism.
-The results are averaged over 3 runs and are consistent between each run.
-The first uncontrolled real Internet setup consists of a residential client in Bari that accesses the network using an ADSL modem. The server is located at the University of Rome, La Sapienza.
-Figure 3. Satellite testbed scenario.
-The second real network testbed is made of users accessing Starlink and connected to other Internet users as in Figure 3.
-The connection was established between nodes at Politecnico di Bari, Italy and Karlstad University, Sweden.
-Since the experimental results are affected by the uncontrolled coexisting network conditions, considered QUIC congestion control algorithms are executed in sequence and the sequence is repeated 40 times.
-Pacing is enabled at the application level at the server for all the experiments to ensure that packets are transmitted at evenly spaced intervals, minimizing burstiness.
-[^3] The pacing interval has been set to 200 $\mu s$.
+The propagation delay $T_{p}$ is set using netem queueing disciplines on the server Ethernet interface, while the bandwidth capacity $C$ in Mbps and the buffer size $B$ in bytes are set using the token bucket filter (TBF) mechanism.
+The results have been averaged over 3 runs and appear consistent between each run.
+Figure 3. Single bottleneck testbed
+The other testbed considered is an uncontrolled real Internet setup consisting of a residential client in Bari that accesses the Internet using an ADSL modem.
+The server is located at the University of Rome, La Sapienza.
+Pacing is enabled at the application level to ensure that packets are transmitted at evenly spaced intervals, minimizing burstiness.  [^3] The pacing interval has been set to 200 $\mu s$.
 5. Experimental results
-This section presents the experimental results that have been run to validate QUIC-DC.
-To the purpose, QUIC-DC has been compared with other state-of-the-art congestion control algorithms, namely Westwood+ (Grieco and Mascolo, 2004), Cubic (Ha et al., 2008), New Reno (Henderson et al., 2012), and BBR version 2  (Cardwell et al., 2019).
-Notice that, even though BBRv3 is available, in this work we consider BBRv2 due to its superior fairness compared to BBRv3 (Zeynali et al., 2024).
-The bottleneck link capacity has been set equal to $C$ Mbps and the bottleneck buffer size has been set equal 0.5, 1, 2, 4 times the bandwidth delay product (BDP) where the bandwidth is the bottleneck capacity $C$ and the delay is the round trip propagation delay $T_{p}=RTT_{min}$.
-In the following, we show results in the cases of single-flow and multiple-flow connections.
+QUIC-DC has been compared with other state-of-the-art congestion control algorithms, namely Westwood+ (Mascolo et al., 2001; Grieco and Mascolo, 2004), Cubic (Ha et al., 2008), New Reno (Henderson et al., 2012), and BBR version 2  (Cardwell et al., 2019).In this work BBRv2 has been considered due to its superior fairness compared to BBRv3 (Zeynali et al., 2024).
+The bottleneck buffer size $B$ has been set equal 0.5, 1, 2, 4 times the bandwidth delay product (BDP), where the bandwidth is the bottleneck capacity $C$ and the delay is the round trip propagation delay $T_{p}=RTT_{min}$.
+We will consider single-flow and multiple-flows scenarios.
 5.1. Emulated network results
 5.1.1. Single Flow
-To understand the dynamics, we begin by examining a single-flow connection transferring a 100 MB file.
-QUIC-DC has been tested using different values for the one-way queuing delay threshold, denoted as $OWQD_{th}$ set to 10%, 20%, 50%, and 80% of the bottleneck buffer size expressed in time.
+To understand the dynamics of QUIC-DC, we begin by examining a single-flow connection transferring a 100 MB file.
+QUIC-DC has been tested using different values for the one-way queueing delay threshold $OWQD_{th}$, which has been set equal to 10%, 20%, 50%, and 80% of the bottleneck buffer size expressed in time.
 Figure 4. Round‑trip‑time (RTT) distribution for a single flow – emulated link, buffer=4×BDP
-Table 1 shows the performance metrics of a single flow connections for the case of a buffer size set equal to 4 BDP, where $C$=10 Mbps and $RTT_{min}=T_{p}=$50 ms.
-In this case, QUIC-DC reaches maximum goodput (Gput in the table) with minimum loss percentage compared to the other congestion control algorithms independently of the considered queuing delay threshold $OWQD_{th}$ and lower RTT.
-In comparison to QUIC Cubic, the loss percentage is reduced by 2 orders of magnitude!
+Table 1 shows the performance metrics of a single flow connections for the case of a buffer size set equal to 4 BDP, where $C$=25 Mbps and $RTT_{min}=T_{p}=$50 ms.
+In this case, all the congestion control algorithms achieve comprable goodputs but with larger loss rate in the case of Cubic and BBRv2.
+In comparison to QUIC Cubic, the loss percentage is reduced up to two orders of magnitude.
 Figure 4 shows the corresponding CDF of the measured RTT for all the algorithms under analysis.
@@
 Similar considerations can be drawn from Table 2, which corresponds to the case of bottleneck buffer size set to 2 BDP.
-A nice understanding of the proposed approach is provided by Figure 5, which compares the $RTT$ and $cwnd$ dynamics of QUIC Westwood+ (blue line) and QUIC-DC (red line) in the case of a bottleneck buffer size set to 2 BDP.
+A nice understanding of the meaning of the proposed approach is provided by Figure 5, which compares the $RTT$ and $cwnd$ dynamics of QUIC Westwood+ (blue line) and QUIC-DC (red line) in the case of a bottleneck buffer size set to 2 BDP.
 It can be observed that, while QUIC Westwood+ resets the congestion window when the maximum RTT is reached (dashed green line), QUIC-DC resets the congestion window when the $OWQD$ reaches the delay control threshold set at 80% of the bottleneck buffer size in time.
@@
 Consequently, the observed packet loss ratio was reduced by one order of magnitude (see Table 1).
-When the buffer size is set to 1 BDP or 0.5 BDP, results reported in Table 3 and 4 show that QUIC-DC tends to be less efficient, in terms of goodput.
-It is worth noting that, in the case of congestion control algorithms that experience higher packet loss—such as BBR2—the measurement of packet RTT does not account for retransmitted packets (Karn and Partridge, 1987).
-As a result, the RTT values tend to underestimate the actual delays.
+When the buffer size is set to 1 BDP or 0.5 BDP, the results in Table 3 and Table 4 indicate that QUIC-DC tends to be less efficient in terms of goodput.
+Notably, in this scenario with an undersized bottleneck buffer, both BBRv2 and Cubic achieve higher goodput, with Cubic also exhibiting the lowest loss rate.
+As will be shown in the following, this result disappears in the case of a multi-flow scenario.
 Figure 5. Time evolution of RTT and congestion window – buffer=2×BDP.
@@
 New Reno 7.39 7.42 0.37 52.39 1.10
-Table 4. Single-flow emulation; buffer = $0.5\times\text{BDP}$ ( $C=10\,\text{Mbps}$, $RTT_{\min}=50\,\text{ms}$ ).
+Table 4. Single-flow emulated link; buffer = $0.5\times\text{BDP}$ ( $C=10\,\text{Mbps}$, $RTT_{\min}=50\,\text{ms}$ ).
 CCA Gput Tput Loss RTTavg RTTstd
@@
 5.1.2. Multiple Flows
-Let us now consider the multi-flow scenario shown in Figure 3, where four connections of the same type share the same channel and are initiated at different start times, i.e.
-they enter the channel every 10 seconds, and they leave it in a staggered way as well.
-In this case, the control delay threshold $OWQD_{th}$ is set to 80% for QUIC-DC, and the bottleneck buffer size is set to 2 BDP.
-Table 5. Four concurrent flows – emulated link, buffer=2×BDP. (Average $\pm$ standard deviation).
+The multi-flow scenario shown in Figure 3 is considered.
+In particular, four connections of the same type are started at different times to avoid synchronization.
+In this case, the control delay threshold $OWQD_{th}=80\%$ and the bottleneck buffer size is set to 2 BDP.
+We have measured and averaged goodputs, throughputs, losses and RTTs for each connection.
+Table 5. Four concurrent flows, emulated link – buffer=2×BDP. (Average $\pm$ standard deviation).
 CCA RTTavg Goodput Loss
@@
 As shown, QUIC-DC exhibits the lowest RTT and loss percentage, while maintaining a goodput level comparable to the other considered algorithms.
-It is worth noting that, since the link capacity is set to 10 Mbps, a goodput around 2.5 Mbps indicates that the link is fairly shared among the four connections.
+It is worth noting that, since the link capacity is set to 25 Mbps, a goodput around 6 Mbps indicates that the link is fully utilized and fairly shared among the four connectioncs.
 Figure 6. RTT distribution for four concurrent flows – emulated link, buffer=2×BDP.
 Figure 6 shows the CDF of the RTT in the case of multiple connections over a shared bottleneck.
-The figure shows that QUIC-DC provides lower RTT values compared to the other considered congestion control algorithms.
-Due to space constraints, we are unable to extend our investigation to more complex multiflows scenarios, nor can we explore the friendliness between different types of control algorithms, which are left for further investigation.
+It shows that QUIC-DC provides lower RTT values compared to the other considered congestion control algorithms.
+Due to space constraints, we are here unable to investigate the friendliness between different types of control algorithms.
 5.2. Real network results
-In this section we report results obtained in the real network scenario described in Section 4.
-5.2.1. Terrestrial network
+We consider the real network scenario described in Section 4.
 During the experiments, we have measured $RTT{min}=30\ ms$, $RTT{max}=60\ ms$, which means that there is a bottleneck buffer size equal to $1\ BDP$.
 Table 6 shows that QUIC-DC reduces the loss rate by one order of magnitude whereas the goodput is only slightly lower than the other approaches (note that QUIC Westwood+ corresponds to QUIC-DC $(100\%)$).
-Moreover, Figure 7 shows that QUIC-DC also provides shorter RTTs.
-It should be noted that algorithms experiencing higher loss rates provide RTT delays that are underestimated because retransmitted packets are not considered in RTT measurements (Karn and Partridge, 1987).
-Figure 7. RTT distribution on real terrestrial path.
-Table 6. Terrestrial path, single flow.
+Moreover, Figure 7 shows that QUIC-DC also provides shorter RTT.
+However, it should also be noted that algorithms experiencing higher loss rates provide RTT delays that are underestimated because retransmitted packets are not considered in RTT measurements (Karn and Partridge, 1987).
+In other terms, the algorithms with larger loss-rates provide a distribution of RTT that underestimates the effective delay.
+Figure 7. RTT distribution on real network path.
+Table 6. Single flow, real network connections.
 CCA Gput Tput Loss RTTavg RTTstd
@@
 New Reno 33.935 32.959 0.0765 35.92 12.18
-5.2.2. Starlink LEO network
-We consider the satellite scenario depicted in Figure 3.
-During the experiments, we measured $RTT_{min}=45\ \mathrm{ms}$ and $RTT_{max}\simeq 100\ \mathrm{ms}$, implying a queuing delay of circa 55 ms.
-This corresponds to a buffer size of approximately $1\times\mathrm{BDP}$.
-We’ve conducted experiments with a single flow and with multiple concurrent flows with the same methodology shown in Section 5.1.Table 7 shows that BBRv2 obtains a significant larger goodput, a result that is not kept in the case of multi-flows.
-Table 8 shows that BBRv2 is broken in the multi-flow case, where it gets 1/4 of the QUIC Westwood+ goodput with a loss rate 2.4 times higher.
-Figure8 shows the CDF of the RTTs for all the considered algorithms in the case of single flow.
-QUIC CUBIC obtains a slightly better distribution but with a smaller goodput as shown in Table 7.
-Figure 9 shows the CDF of the RTTs for all the considered algorithms in the case of multiple flows.
-In this case QUIC BBRv2 provides the worst distribution even though it provides the worst goodput (only 5 Mbps whereas QUIC Westwood+ provides 19 Mbps).
-QUIC-DC using different thresholds on the one-way delay can slightly improve the RTT distributions.
-However, it should also be noted that algorithms experiencing higher loss rates provide RTT delays that are underestimated because retransmitted packets are not considered in RTT measurements (Karn and Partridge, 1987).
-In other terms, the algorithms with larger loss-rates provide a distribution of RTT that underestimates the effective delay.
-Figure 8. RTT distribution on LEO satellite path – single flow
-Table 7. Starlink LEO satellite path, single flow.
-CCA Gput Tput Loss RTTavg RTTstd
-(Mbps) (Mbps) (%) (ms) (ms)
-QUIC-DC (20%) 4.81 4.82 0.13 60.58 7.51
-QUIC-DC (50%) 18.06 18.08 0.09 62.30 9.41
-Westwood+ 20.12 20.29 0.80 63.21 11.82
-BBRv2 29.85 30.17 1.04 67.70 27.02
-Cubic 15.07 15.19 0.82 59.29 18.31
-New Reno 9.45 9.64 1.96 66.05 20.83
-Figure 9. RTT distribution on Starlink LEO satellite path – four concurrent flows
-Table 8. Starlink LEO satellite path, four concurrent flows.
-CCA RTTavg Goodput Loss
-(ms) (Mbps, JFI) (%)
-QUIC-DC (20%) $66.36\pm 0.35$ $4.05\pm 0.07$ (1.000) $0.10\pm 0.01$
-QUIC-DC (50%) $72.29\pm 0.98$ $8.70\pm 0.61$ (0.996) $0.09\pm 0.02$
-Westwood+ $81.72\pm 1.27$ $19.17\pm 0.95$ (0.998) $1.02\pm 0.45$
-BBRv2 $87.04\pm 10.50$ $5.09\pm 1.60$ (0.931) $2.44\pm 0.63$
-Cubic $74.76\pm 2.78$ $15.28\pm 1.63$ (0.992) $0.81\pm 0.09$
-New Reno $78.70\pm 2.06$ $14.81\pm 1.49$ (0.992) $0.93\pm 0.41$
 6. Conclusions
-In this paper, we presented QUIC-DC, an algorithm that explicitly controls the queuing delay of a communication connection.
-The core idea is to estimate the connection one-way queuing delay in order to trigger an early reaction to congestion.
-The idea has been implemented in QUIC using the TCP Westwood+ congestion control.
-QUIC-DC has been evaluated against QUIC NewReno, QUIC Westwood+, QUIC Cubic, and QUIC BBR2.
+QUIC DC controls the delay of an Internet connection by measuring the one-way queueing delay to trigger an early reaction to congestion.
+The idea has been implemented in QUIC using the TCP Westwood+ congestion control and has been evaluated against QUIC NewReno, QUIC Westwood+, QUIC Cubic, and QUIC BBRv2.
 Experimental results show that the cumulative distribution function of the round-trip time can be shifted to lower values, while the packet loss rate can be reduced by an order of magnitude.
+BBRv2 has shown serious drawbacks when multiple flows share the bottleneck.
+Acknowledgements.
+This work has been funded by the “LOREN” Project,
+n. 20223Y85JN, under the PRIN (Progetti di Ricerca d’Interesse Nazionale) 2022 programme of the Italian Ministry of Research.